### v1

@@ 4.4 Approximate DOS of the RP model
Until now, the approximate expressions for $G_{C}(z)$ we have derived above are valid for any operator $A$ with a well-defined DOS, which is then added with an operator $B$ having a semi-circle density. Now specializing to the RP model considered in this paper, we see that the eigenvalue distribution for the diagonal matrix $A$ is a Gaussian distribution with zero mean and variance $\sigma^{2}=4/N$. Hence, denoting the operator $C$ as the Hamiltonian $H$ of the RP model, we have
@@ 6.2 Operator statistics at different times
At initial time $t=0$, the operator statistics are independent of the Hamiltonian and thus follow the multinomial distribution, depending on their number of distinct eigenvalues (see Table 1). In the ergodic phase, the eigenvalue statistics of the operator exhibit convergence toward the arcsine distribution at the chosen timestamp, as predicted by free probability theory Camargo:2025zxr . This behavior reflects the expected random matrix-like dynamics within the ergodic regime. In contrast, the eigenvalue distributions in the fractal and localized phases deviate significantly from the arcsine distribution at the same timestamp. These deviations highlight the role of different nature of the quantum dynamics in the fractal and localized regimes, and we can speculate that quantum dynamics in these regimes of the Hamiltonian do not drive the operators (here $S_{1}(0)$ and $S_{L}(0)$) towards freeness (see also the discussion on the late time behavior of these statistics below).
@@
A natural question one can now ask is the following: do the eigenvalue statistics in the fractal and localized phases eventually approach the free probability prediction at sufficiently late times? [^13] The specific shape of the distribution might hold its own intrinsic interest; however, our primary focus is on whether it converges to the free probability distribution. We thank Kohei Kawabata and Hal Tasaki for discussions on this point. The results presented in Fig. 9 suggest that they do not. At very late times $t\sim O(10^{4})$, for all the spins, the operator eigenvalue distribution in the fractal phase stabilizes into a distinct form, which remains fundamentally different from the free probability prediction. However, there are several spikes observed, exactly at the same point, keeping the memory of the initial distribution at $t=0$. This indicates that the fractal phase sustains its intermediate, non-ergodic characteristics even at prolonged timescales.
@@ 6.3 Analysis of the free time
We apply the KS test to assess whether the empirical distributions for the eigenvalues of $A(0)+A(t)$ follow the prediction from free probability. In Fig. 12, we show the $p$-value of the test as a function of time for different values of $\gamma$. The figure shows that at early times the $p$-value takes values smaller than 0.05, indicating that the empirical distribution and the free probability prediction are not the same. However, as time increases, the $p$-value increases, taking values larger than 0.05, indicating that the free probability prediction has emerged. The free time is computed as the time at which the $p$-value crosses the value 0.05.161616We also checked with other tests such as Pearson’s $\chi^{2}$ test and confirm the strikingly similar behavior of free time, as in Fig. 11.
@@ 7 Conclusion and outlook
Within the framework of free probability, we revisit the calculation of the DOS for the Rosenzweig–Porter (RP) random matrix ensemble and study the spectral statistics of operators, highlighting the approximate emergence of asymptotic freeness in this model.
@@
Since the RP model is defined as the sum of two independent random matrices, the two terms are free in the sense of free probability in the limit when the dimension of the matrices is sufficiently large. This enables the use of free probability techniques to compute the DOS. Using a general approach based on subordination functions, we derive an explicit expression for the DOS of the RP model, reproducing previous results obtained via the replica technique Venturelli:2022hka . Our analysis focuses on the fractal phase, where we express the DOS as perturbative corrections to the Gaussian profile characteristic of the localized phase. As illustrated in Fig. 3, the DOS evolves smoothly from a Wigner semicircle to a Gaussian distribution as the parameter $\gamma$ increases. In particular, during the transition from the ergodic to the fractal phase, the DOS resembles a perturbed Wigner semicircle. It would be interesting to extend the methodology developed here to compute the DOS in this intermediate regime.
@@
It has been recently proposed that chaotic dynamics drives the components of a sum of two operators, one of which is time-evolved, toward asymptotic freeness with respect to non-evolved ones Camargo:2025zxr . However, the precise mechanisms underlying the emergence of freeness remain unclear. For instance, is it sufficient for the spectral statistics of the Hamiltonian to follow random matrix theory? What role do the eigenvectors play in this process? To address these questions, we studied the emergence of free probability predictions for the spectrum of operators of the form $S_{1}+S_{L}(t)$ in the RP model, an ensemble of random matrix Hamiltonians where the rotational invariance of the classical Gaussian ensembles is broken explicitly by introducing a parameter, and depending on the value of this parameter the energy eigenvectors show different types of localization properties. Our analysis indicates that such behavior appears only in the ergodic phase; in contrast, free probability fails to accurately describe the operator spectrum in the fractal and localized phases, as expected.
@@
Interestingly, in the fractal phase, the system exhibits partial signatures of freeness, yet retains memory of the spectrum at $t=0$. This suggests that the emergence of freeness is not solely governed by spectral statistics, but also by the structure of the eigenvectors. While the fractal phase still displays level repulsion—as reflected in the average $r$-parameter—the corresponding eigenvectors are not fully delocalized. This points to a natural conclusion: the presence of level repulsion in the Hamiltonian alone is not sufficient for the emergence of freeness in operator statistics; extended, non-localized eigenvectors are also essential. It would be interesting to explore this direction further and test its validity in other models that exhibit fractal or multifractal phases Khaymovich:2021tkj .
@@ Acknowledgements
We thank Yichao Fu, Kohei Kawabata, Tokiro Numasawa, Tanay Pathak, Hal Tasaki, and Masafumi Udagawa for fruitful discussions. We also thank Jonah Kudler-Flam for useful comments on the draft. P.N. acknowledges the hospitality of Gwangju Institute of Science and Technology (GIST), Gakushuin University and the Institute of Solid State Physics (ISSP), University of Tokyo, during the final stages of the work. This work was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT & Future Planning (NRF-2021R1A2C1006791, NRF-2020R1I1A2054376), the framework of international cooperation program managed by the NRF of Korea (RS-2023-NR119907, RS-2024-00440725), the Creation of the Quantum Information Science R&D Ecosystem (Grant No. 2022M3H3A106307411) through the NRF funded by the Korean government (Ministry of Science and ICT), the Gwangju Institute of Science and Technology (GIST) research fund (Future leading Specialized Resarch Project, 2025) and the Al-based GIST Research Scientist Project grant funded by the GIST in 2025. V.J. and H.C. are supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2022R1I1A1A01070589, and RS-2023-00248186). The work of P.N. is supported by the JSPS Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe” No. 21H05190.
@@ Appendix B Free additive convolution of spin-$s$ operators
$\rho_{s}(\lambda)=\frac{1}{2s+1}\sum_{j=-s}^{s}\delta(\lambda-s)\,.$ (65)
@@
$G_{s}(z)=\int\frac{\rho_{s}(\lambda)}{z-\lambda}=\frac{1}{2s+1}\sum_{j=-s}^{s}\frac{1}{z-s}\,.$ (66)

### v2

@@ 4.4 Approximate DOS of the RP model
Until now, the approximate expressions for $G_{C}(z)$ we have derived above are valid for any operator $P$ with a well-defined DOS, which is then added with an operator $Q$ having a semicircle density. Now specializing to the case of the RP model considered in this paper, we see that the eigenvalue distribution for the diagonal matrix $P\equiv A$ is a Gaussian distribution with zero mean and variance $\sigma^{2}=4/N$, and $Q\equiv B$. Hence, denoting the operator $C$ as the Hamiltonian $H$ of the RP model, we have
@@ 5.2 Approximate analytical expressions for the two-point function
5.3 Cumulative OTOCs
As observed in the previous section, higher $n$-point functions exhibit a fast decay in the ergodic regime. Although this behavior persists in the fractal phase, such decay is significantly slower compared with the ergodic phase. Conversely, in the localized regime, these functions seem to retain their initial values and remain non-decaying for large time scales. Given that the decay of all $n$-point functions signals the emergence of freeness, we can characterize this emergence by considering the sum of such $n$-point functions. Such a sum would characterize the cumulative correlation between the operators. If $f(n)$ is a positive function of $n\in\mathbb{N}$ and $g(x)$ is a positive and continuous function of $x\in\mathbb{R}$, then we consider the partial sum of $p$ weighted and positive combinations of $n$-point functions $\mathcal{F}_{p}^{(\gamma)}(f,g\,;t)$ given by
$\mathcal{F}_{p}^{(\gamma)}(f,g\,;t):=\sum_{n=1}^{p}f(n)\,g(F_{2n}^{(\gamma)}(t))\,,$ (55)
where $F_{2n}^{(\gamma)}(t)$ are given by (50). Here, the role of $f(n)$ is to provide a specific weight to each of the functions $g(F_{2n}^{(\gamma)}(t))$ in the partial sum, while $p$ represents the maximum number of $n$-point functions considered, and taking the limit $p\rightarrow\infty$ incorporates all sums of higher $n$-point functions. Note that since $g$ is a positive function, $\mathcal{F}_{p}^{(\gamma)}(f,g\,;t)$ will be zero if and only if the individual $g(F_{2n}^{(\gamma)}(t))$ vanish. Thus, this allows us to characterize the vanishing of all $n$-point functions using a single quantity.
The simplest positive function $g$ that we can consider is $g(x)=x^{2}$, leading to
$\mathcal{F}_{p}^{(\gamma)}(f\,;t):=\sum_{n=1}^{p}f(n)(F_{2n}^{(\gamma)}(t))^{2}=\sum_{n=1}^{p}f(n)\langle(S_{1}(0)S_{1}(t))^{n}\rangle^{2}\,.$ (56)
The square is in a similar spirit to Vallini and Pappalardi Vallini:2024bwp , where $g(x)=\log|x|$, without the cumulative sum. On the other hand, the sum in (56) thus corresponds to the cumulative contribution of weighted $n$-point squared OTOCs. At the same time, there is a large degree of freedom in choosing the weight $f(n)$. One way to choose it would be to fix the limiting value of the cumulative sum at $t=0$. In such a case, and for spin operators in the $p\rightarrow\infty$ limit, we have
$\mathcal{F}_{\infty}^{(\gamma)}(f\,;0):=\sum_{n=1}^{\infty}f(n)\,,$ (57)
since $\langle(S_{1}(0)S_{1}(0))^{n}\rangle^{2}=1$. One criterion for choosing $f(n)$ would be to ensure the convergence of  (57). Choosing, for example, $f(n)=n^{-k}$ for $\mathbb{Z}\ni k>1$ leads to $\mathcal{F}_{\infty}^{(\gamma)}(0)=\zeta(k)$, where $\zeta(k)$ is the Riemann zeta function. Choosing $f(n)=n^{-1}$ also provides an alternative weight, which gives a finite result for finite $p$ (the Harmonic number $H_{p}$) but which otherwise diverges in the $p\rightarrow\infty$ limit. [^14] We remark that such a choice is related to the convention of an independent study of operator spreading using free probability theory Shreyatalk .
Figures 6 and 7 illustrate the behaviour of the cumulative partial sum $\mathcal{F}_{p}^{(\gamma)}(t):=\mathcal{F}_{p}^{(\gamma)}(f\,;t)$ for the choice of weights $f(n)=1/n$ and $f(n)=1/n^{2}$ respectively in the ergodic (left) and fractal (right) regimes, for $s=1/2$ and $L=8$. Since higher-point functions decay faster than lower-point functions Cotler:2017jue , in practice, only a finite $p$ suffices to produce reasonable results.
Figure 6: Left: The cumulative OTOC (56) for the ergodic phase, keeping up to $p$-th order. Here we choose $f(n)=1/n$ as the weight factor in (56). The black dashed curve is the exponential decay obtained by summing up an infinite number of terms, given by Eq. (59). Right: The behavior of OTOC sum in the fractal phase, with an inset showing the late-time behavior. The parameters are $s=1/2$, $L=8$ and $50$ ensemble averages of RP Hamiltonian is considered.
Figure 7: Left: The cumulative OTOC (56) for the ergodic phase, keeping up to $p$-th order considering $f(n)=1/n^{2}$. The saturation for the infinite sum is given by $\zeta(2)=\pi^{2}/6$. Right: The behavior of OTOC sum in the fractal phase, with an inset showing the late-time behavior. The parameters are $s=1/2$, $L=8$ and $50$ ensemble averages of RP Hamiltonian is considered.
Assuming an exponential decay of the $2$-point OTOCs Vallini:2024bwp , the cumulative partial sum for a weight of the form $f(n)=n^{-k}$ for $\mathbb{Z}\ni k>0$ becomes
$\sum_{n=1}^{p}\frac{1}{n^{k}}e^{-2n\chi t}=-e^{-2(p+1)\chi t}\Phi(e^{-2\chi t},k,p+1)+\textrm{Li}_{k}(e^{-2\chi t})\,,$ (58)
[…]
$\sum_{n=1}^{\infty}\frac{1}{n^{k}}e^{-2n\chi t}=\textrm{Li}_{k}(e^{-2\chi t})\approx e^{-2\chi t}\,\quad(\mathrm{large}~{}t)\,.$ (59)
This is indeed representative of the numerical results for the ergodic phase (left) in Figs. 6 and 7, where the horizontal red dashed line in the latter case corresponds precisely to $\zeta(2)=\pi^{2}/6$. The exponential decay is represented by the black dashed line in Figs. 6 and 7 (left), demonstrating the eventual vanishing of the cumulative OTOC. The $\gamma=0$ result primarily shows the result when the Hamiltonian is strictly controlled by the GOE matrices in the RP Hamiltonian. The same conclusion holds for other classes of Gaussian ensembles.
As an additional remark, it should also be noted that if we instead assume a slower decay, for example, a power-law behaviour of the form $(F_{2n}^{(\gamma)}(t))^{2}=t^{-2\chi n}$, then the cumulative sum for $p\rightarrow\infty$ would lead to $\mathcal{F}_{\infty}^{(\gamma)}(t)\approx t^{-2\chi}$ in the large time regime. This is consistent with findings of the decay of OTOCs and higher-point OTOCs in random matrix models Cotler:2017jue .
In the fractal phase, the cumulative sum undergoes a decay, albeit at a much slower rate compared to the ergodic regime. As illustrated in the inset of Figs. 6 and 7 (right), at late times, the cumulative OTOC does not approach zero; rather, it decays to a small, non-zero value. This phenomenon reflects a form of “memory” retained within the operator statistics, which we will see in the subsequent section. In contrast, within the localized phase, the cumulative partial sum exhibits minimal decay, remaining almost constant at its initial value.
@@ 6.2 Operator statistics at different times
At initial time $t=0$, the operator statistics are independent of the Hamiltonian and thus follow the multinomial distribution, depending on their number of distinct eigenvalues (see Table 1). This is also known as classical convolution. In the ergodic phase, the eigenvalue statistics of the operator exhibit convergence toward the arcsine distribution at the chosen timestamp, as predicted by free probability theory Camargo:2025zxr . This behavior reflects the expected random matrix-like dynamics within the ergodic regime. In contrast, the eigenvalue distributions in the fractal and localized phases deviate significantly from the arcsine distribution at the same timestamp. These deviations highlight the role of different nature of the quantum dynamics in the fractal and localized regimes, and we can speculate that quantum dynamics in these regimes of the Hamiltonian do not drive the operators (here $S_{1}(0)$ and $S_{L}(0)$) towards freeness (see also the discussion on the late time behavior of these statistics below).
@@
A natural question one can now ask is the following: do the eigenvalue statistics in the fractal and localized phases eventually approach the free probability prediction at sufficiently late times? [^15] The specific shape of the distribution might hold its own intrinsic interest; however, our primary focus is on whether it converges to the free probability distribution. We thank Kohei Kawabata and Hal Tasaki for discussions on this point. The results presented in Fig. 11 suggest that they do not. At very late times $t\sim O(10^{4})$, for all the spins, the operator eigenvalue distribution in the fractal phase stabilizes into a distinct form, which remains fundamentally different from the free probability prediction. However, there are several spikes observed, exactly at the same point, keeping a memory of the initial distribution at $t=0$. This indicates that the fractal phase sustains its intermediate, non-ergodic characteristics even at prolonged timescales. While we have considered two separate operators in the operator statistics, this effect of memory arises due to the non-vanishing contribution of the cumulative OTOC, as we observed in Fig. 6 (right) and Fig. 7 (right).
@@
In summary, we conclude that in the ergodic phase, the classical convolution of two commuting operators evolves into the free convolution at late times. In contrast, in the fractal and localized phases, this transition does not occur.
@@ 6.3 Analysis of the free time
We apply the KS test to assess whether the empirical distributions for the eigenvalues of $A(0)+A(t)$ follow the prediction from free probability. In Fig. 14, we show the $p$-value of the test as a function of time for different values of $\gamma$. The figure shows that at early times the $p$-value takes values smaller than 0.05, indicating that the empirical distribution and the free probability prediction are not the same. However, as time increases, the $p$-value increases, taking values larger than 0.05, indicating that the free probability prediction has emerged. The free time is computed as the time at which the $p$-value crosses the value 0.05.181818We also checked with other tests such as Pearson’s $\chi^{2}$ test and confirm the strikingly similar behavior of free time, as in Fig. 13.
@@ 7 Conclusion and outlook
Within the framework of free probability, we revisit the calculation of the DOS for the Rosenzweig–Porter (RP) random matrix ensemble and study the spectral statistics of operators, highlighting the approximate emergence of asymptotic freeness in this model. Since the RP model is defined as the sum of two independent random matrices, the two terms are free in the sense of free probability in the limit when the dimension of the matrices is sufficiently large. This enables the use of free probability techniques to compute the DOS. Using a general approach based on subordination functions, we derive an explicit expression for the DOS of the RP model, reproducing previous results obtained via the replica technique Venturelli:2022hka . Our analysis focuses on the fractal phase, where we express the DOS as perturbative corrections to the Gaussian profile characteristic of the localized phase. As illustrated in Fig. 3, the DOS evolves smoothly from a Wigner semicircle to a Gaussian distribution as the parameter $\gamma$ increases. In particular, during the transition from the ergodic to the fractal phase, the DOS resembles a perturbed Wigner semicircle. It would be interesting to extend the methodology developed here to compute the DOS in this intermediate regime.
@@
It has been recently proposed that chaotic dynamics drives the components of a sum of two operators, one of which is time-evolved, toward asymptotic freeness with respect to non-evolved ones Camargo:2025zxr . However, the precise mechanisms underlying the emergence of freeness remain unclear. For instance, is it sufficient for the spectral statistics of the Hamiltonian to follow random matrix theory? What role do the eigenvectors play in this process? To address these questions, we studied the emergence of free probability predictions for the spectrum of operators of the form $S_{1}+S_{L}(t)$ in the RP model, an ensemble of random matrix Hamiltonians where the rotational invariance of the classical Gaussian ensembles is broken explicitly by introducing a parameter, and depending on the value of this parameter the energy eigenvectors show different types of localization properties. Our analysis reveals that the transition from classical convolution to free convolution manifests exclusively in the ergodic phase. In contrast, neither the fractal phase nor the localized phase exhibits this free convolution behavior.
@@
Interestingly, in the fractal phase, the system exhibits partial signatures of freeness, yet retains memory of the spectrum at $t=0$. This memory is associated with the non-vanishing cumulative OTOCs at late times, which suggests that the emergence of freeness is not solely governed by spectral statistics, but also by the structure of the eigenvectors. While the fractal phase still displays level repulsion—as reflected in the average $r$-parameter—the corresponding eigenvectors are not fully delocalized. This points to a natural conclusion: the presence of level repulsion in the Hamiltonian alone is not sufficient for the emergence of freeness in operator statistics; extended, non-localized eigenvectors are also essential. It would be interesting to explore this direction further and test its validity in other models that exhibit fractal or multifractal phases Khaymovich:2021tkj .
@@ Acknowledgements
We thank Yichao Fu, Koji Hashimoto, Kohei Kawabata, Tokiro Numasawa, Silvia Pappalardi, Tanay Pathak, Hal Tasaki, and Masafumi Udagawa for fruitful discussions. We also thank Jonah Kudler-Flam for useful comments on the draft. P.N. acknowledges the hospitality of Gwangju Institute of Science and Technology (GIST), Gakushuin University and the Institute of Solid State Physics (ISSP), University of Tokyo, during the final stages of the work. This work was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT & Future Planning (NRF-2021R1A2C1006791, NRF-2020R1I1A2054376), the framework of international cooperation program managed by the NRF of Korea (RS-2023-NR119907, RS-2024-00440725), the Creation of the Quantum Information Science R&D Ecosystem (Grant No. 2022M3H3A106307411) through the NRF funded by the Korean government (Ministry of Science and ICT), the Gwangju Institute of Science and Technology (GIST) research fund (Future leading Specialized Resarch Project, 2025) and the Al-based GIST Research Scientist Project grant funded by the GIST in 2025. V.J. and H.C. are supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2022R1I1A1A01070589, and RS-2023-00248186). The work of P.N. is supported by the JSPS Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe” No. 21H05190.
@@ Appendix B Free additive convolution of spin-$s$ operators
$\rho_{s}(\lambda)=\frac{1}{2s+1}\sum_{j=-s}^{s}\delta(\lambda-j)\,.$ (74)
@@
$G_{s}(z)=\int\frac{\rho_{s}(\lambda)}{z-\lambda}=\frac{1}{2s+1}\sum_{j=-s}^{s}\frac{1}{z-j}\,.$ (75)